OpenStudy (anonymous):

1+3 solve that please

OpenStudy (uri):

4..

OpenStudy (anonymous):

4 IN MATHMATICAL TERMS.

OpenStudy (austinl):

1600

OpenStudy (anonymous):



OpenStudy (anonymous):

\(1+3=1+1+1+1=0.5+0.5+0.5+0.5\) Do it be hand and you get \(4\)